OpenStudy (debbieg):

How long will it take money to double if it is invested at 5% compounded annually? ..... I am pretty sure I have the right answer but the back of the text says something else. Just want to confirm that I'm right and it's wrong! lol

OpenStudy (debbieg):

I get 14.2 years. 2=(1+0.05)^t ln2=t*ln(1.05) t=ln2/(ln1.05)=14.2 Am I missing something? Book says 15 years.

OpenStudy (anonymous):

i don't think you're wrong. my ideas: -maybe they are rounding up since it is unlikely to be 15.000 years. -unlikely: maybe they used the "rule of 72" but 72/5 = 14.4. still not justifiable to round up. but this is how this problem is meant to be done

OpenStudy (debbieg):

Thanks. Certainly not the first time an answer in the back of the book is wrong, but I wanted to be sure I wasn't making some dumb mistake. I don't think it's rounding, because the other part of the same problem is to find the time under continuous compounding, and there they have 13.9 (which I get too).

OpenStudy (anonymous):

lol i hate when that happens :P i've wasted so much time with wrong textbook answers